# Condition-Aware Multimodal Fusion for Robust Semantic Perception of Driving Scenes

### v1

@@ Abstract
Leveraging multiple sensors is crucial for robust semantic perception in autonomous driving, as each sensor type has complementary strengths and weaknesses. However, existing sensor fusion methods often treat sensors uniformly across all conditions, leading to suboptimal performance. By contrast, we propose a novel, condition-aware multimodal fusion approach for robust semantic perception of driving scenes. Our method, CAFuser, uses an RGB camera input to classify environmental conditions and generate a Condition Token that guides the fusion of multiple sensor modalities. We further newly introduce modality-specific feature adapters to align diverse sensor inputs into a shared latent space, enabling efficient integration with a single and shared pre-trained backbone. By dynamically adapting sensor fusion based on the actual condition, our model significantly improves robustness and accuracy, especially in adverse-condition scenarios. We set the new state of the art with CAFuser on the MUSES dataset with 59.7 PQ for multimodal panoptic segmentation and 78.2 mIoU for semantic segmentation, ranking first on the public benchmarks.
@@ I Introduction
Extensive experiments show that our method, CAFuser (Condition-Aware Fuser), effectively learns to weigh different modalities in the fusion according to the present condition. CAFuser sets the new state of the art in multimodal panoptic and semantic segmentation, ranking first on the public MUSES [3] benchmarks.
@@ III Method
We build on the multimodal fusion approach introduced with the MUSES [3] dataset, which integrates multiple sensor modalities for panoptic segmentation by adapting Mask2Former [14] with a multi-branch multimodal backbone and an MWCA fusion block [6] for each feature level. As OneFormer [17] has surpassed Mask2Former, we adapt the former multi-branch fusion architecture from MUSES and integrate it into the OneFormer segmentation framework, which we call “OneFormer-MUSES” and use it as a strong multimodal baseline. Unless otherwise stated, we use the Swin-T [37] backbone for all experiments due to computational constraints.
@@ III-A Multimodal Adapter
As displayed in Fig. 2, we use a shared backbone for all modalities instead of individual backbones for each modality, to simplify and optimize the fusion process. We initialize the backbone with ImageNet pre-training, which corresponds to the RGB camera modality. For pre-processing, we follow MUSES and project each sensor’s data (e.g., lidar, radar) as a 3-channel image onto the RGB plane. Further, we normalize each modality over the entire dataset to ensure an input representation consistent with the RGB modality.
To significantly reduce model parameters while maintaining competitive performance (cf. Sec. IV-B), we employ a lightweight adapter that consists of a 2-layer MLP with a 4x reduction in hidden dimensions [5], enabling efficient feature transformation. A learnable parameter $\alpha$ controls the weighting of adapted and original features. Each modality and each feature map from the Swin backbone’s 4-level feature pyramid is adapted using an individual adapter at each stage. For 4 modalities and 4 feature levels this results in 16 individual lightweight adapters. In Sec. IV-B we show that this setup allows us to reduce the parameters by 54% without any loss in performance.
@@ IV Experiments
We carry out our experiments and extensively ablate our components on the MUSES dataset [3], which provides 2500 real-world multimodal driving scenes with high-quality panoptic segmentation annotations, making it a crucial benchmark for assessing model performance in adverse weather conditions. The dataset includes 8 different weather conditions, combining clear, rain, snow, and fog with both day and nighttime, allowing to test the ability of a segmentation network to parse highly degraded inputs. Each scene includes a standard RGB camera, lidar, radar, event camera, and metadata describing scenes’ weather conditions. We use the MUSES SDK for projecting each modality onto the RGB image plane and we implement our model on the Detectron2 [41] framework. As a backbone, we use Swin tiny (T) [37]. CAFuser is trained using the standard training setup of OneFormer [17], with 960 epochs and a batch size of 8. Following [3], we use a 20% random drop rate for each modality.
@@ IV-A Comparison to The State of The Art
In Table I, we compare our model with existing panoptic segmentation methods, both using RGB-camera-only inputs and multimodal inputs. For RGB-camera-only panoptic segmentation, we compare to strong baselines such as Mask2Former [14], MaskDINO [16], and OneFormer [17]. Among these methods, OneFormer achieves the highest performance, with a PQ score of 55.2%, highlighting the strength of this architecture for RGB-camera-based segmentation. For multimodal panoptic segmentation, OneFormer-MUSES [17, 3] sets a strong baseline. However, our method, which introduces CAF, achieves a new state-of-the-art result with a PQ score of 59.7%, surpassing all previous methods, with only half the parameters of OneFormer-MUSES.
As multimodal panoptic segmentation is still an emerging field, we further benchmark our model against the state of the art in multimodal semantic segmentation. Given the OneFormer head simultaneously solves panoptic and semantic segmentation, we obtain one CAFuser model solving both tasks. Table II highlights the performance of our model in comparison to top methods such as CMNeXt [7] and GeminiFusion [8]. While these multimodal methods do improve over the RGB-camera-only methods, our CAFuser-CA2 model significantly outperforms all multimodal and RGB-only methods and sets the new SOTA. This validates the strength of CAF, which allows optimal sensor integration under diverse environmental conditions and thus superior performance both in panoptic and semantic segmentation.
TABLE I: Comparison of panoptic segmentation methods on the MUSES test set. : multimodal OneFormer according to Sec. III, C: RGB Camera, L: Lidar, R: Radar, E: Event Camera.
Method Modalities Backbone PQ $\uparrow$
Mask2Former [14] C Swin-T 46.9
MaskDINO [16] C Swin-T 49.2
OneFormer [17] C Swin-T 55.2
HRFuser [6] CLRE HRFuser-T 43.5
MUSES [3] CLRE 4xSwin-T 53.9
OneFormer-MUSES [17] CLRE 4xSwin-T 59.3
CAFuser-CAA CLRE Swin-T 59.4
CAFuser-CA2 (Ours) CLRE Swin-T 59.7
@@ IV-B Ablation Studies
Feature Adapter: We ablate the effect of our proposed feature adapter and CAF module in Tab. III. Using a shared backbone (row 3) significantly reduces the number of parameters (-55%) compared with our baseline (row 2), but also results in a significant performance drop (-0.9% PQ). Adding our proposed feature adapter (row 4) gains back all the lost performance (+0.9% PQ) while still having less than half of the original parameters. Adding our CAF module (row 5) gains another 0.4% in PQ, surpassing the performance of the larger, baseline model.
TABLE III: Ablation study of our proposed modules. Using a shared backbone reduces the parameters by 55% while the adapter and CAF module increase the PQ to a new SOTA. : multimodal OneFormer according to Sec. III
@@
CAA Fusion: We conduct an ablation study to evaluate the impact of our CAA fusion method. As shown in Table IV, we compare four fusion strategies: mean, random weights, learned weights, and CAA Fusion, which dynamically predicts modality weights using the CT. For the basic method that takes a simple mean over all modalities, we can see a large improvement of +1% PQ when adding the adapter (row 2) to the no-adapter baseline, highlighting the benefit of adapting features before fusion. Random fusion, where weights are randomly assigned to each modality, underperformed with a PQ of 58.5%. Learning static weights for each modality, without considering environmental conditions, reaches a similar performance to the mean fusion. Finally, our proposed CAA Fusion, which dynamically adjusts weights based on the CT, outperforms all other strategies with a PQ of 59.4%, demonstrating the effectiveness of adapting the fusion process to environmental conditions.
TABLE IV: Ablation study using all four modalities on different weighted fusion strategies for our CAA module.
@@
CA2 Fusion: In Tab. V, we investigate the detailed design of our CA2 fusion and how to best utilize the CT for CAF. Specifically, we examine whether the CT should be appended to the RGB queries (Q), to the keys and values (K&V) of the secondary modalities, or to both. Our proposed approach (row 3), where the CT is only appended to the RGB queries, achieves the highest PQ score of 59.7%. Appending the CT to the secondary modalities’ keys and values yields a slight decrease in performance. This motivates our design choice of appending the CT to the queries to help guide cross-attention and follows our intuition from Sec. III-B.
TABLE V: Ablation study on CA2 fusion design. The Condition Token is appended to the RGB queries (Q) or the secondary modalities’ keys and values (K&V).
@@
Condition Loss: In Table VI, we assess the impact of our contrastive condition loss, by training an identical network, with and without applying the loss on the CT. As this loss is designed to ensure the CT effectively captures environmental conditions, we want to investigate if the model could learn this internally without additional supervision. Adding the condition loss improves the performance by +0.4% in PQ. This demonstrates that our contrastive loss is essential as an additional supervision to efficiently guide CAF.
TABLE VI: Ablation study on the guidance of the verbo-visual contrastive loss between $Q_{text}$ and the Condition token.
@@
Figure 6: Qualitative segmentation results on MUSES with visualization of the four input modalities. Best viewed on a screen at full zoom.
Condition Encoding: In Table VII, we ablate the use of different condition prompt encodings in the CT generation. Using a high-level condition description, such as simply classifying weather with a single token as e.g. clear or foggy, yields the same performance as without CAF (PQ 59.2%). However, our proposed detailed condition prompt construction, which captures nuanced aspects of the environment (e.g., distinguishing between snow on the ground versus snow in the air), results in a significant improvement of +0.5% in PQ, achieving a PQ of 59.7%. This suggests that detailed condition prompts provide more fine-grained information for the fusion process, enabling more effective sensor integration based on the actual environmental context.
Modalities: Table VIII shows the performance of our model with different combinations of input modalities. Starting with RGB-only input (55.7% PQ), we observe consistent performance gains with each added modality: +3.0% PQ for lidar, +0.6% PQ for radar, and an additional +0.4% PQ for event camera, reaching 59.7% PQ. These results demonstrate that our approach effectively leverages all modalities, with each additional sensor contributing to performance. The flexibility of our design allows to generalize well across various input combinations.
TABLE VII: Ablation study on the condition prompt encodings. 1: simple high-level condition prompt (e.g. [“clear”]) 2: detailed condition prompts according to Sec. III-B.
@@
TABLE VIII: Ablation for CAFuser wit different input modalities. Each added modality improves the PQ.
@@ V Conclusion
In this work, we introduced CAFuser, a novel condition-aware multimodal fusion framework for robust semantic perception in autonomous driving. By employing a shared backbone and modality-specific feature adapters, CAFuser efficiently aligns diverse sensor inputs into a common latent space while significantly reducing the model complexity. Our attention-based condition-aware fusion module dynamically adapts to environmental conditions guided by a Condition Token that is learned from the RGB input. This dynamic fusion enhances robustness and accuracy under challenging weather scenarios. We demonstrated that our method outperforms competing approaches both in panoptic and semantic segmentation, setting the new state of the art on the MUSES dataset. Extensive ablation studies demonstrate the effectiveness of our proposed modules. These advancements make CAFuser a promising solution for enhancing semantic perception in autonomous driving and robotics, particularly under adverse environmental conditions.

### v2

# CAFuser: Condition-Aware Multimodal Fusion for Robust Semantic Perception of Driving Scenes

@@ Abstract
Leveraging multiple sensors is crucial for robust semantic perception in autonomous driving, as each sensor type has complementary strengths and weaknesses. However, existing sensor fusion methods often treat sensors uniformly across all conditions, leading to suboptimal performance. By contrast, we propose a novel, condition-aware multimodal fusion approach for robust semantic perception of driving scenes. Our method, CAFuser, uses an RGB camera input to classify environmental conditions and generate a Condition Token that guides the fusion of multiple sensor modalities. We further newly introduce modality-specific feature adapters to align diverse sensor inputs into a shared latent space, enabling efficient integration with a single and shared pre-trained backbone. By dynamically adapting sensor fusion based on the actual condition, our model significantly improves robustness and accuracy, especially in adverse-condition scenarios. CAFuser ranks first on the public MUSES benchmarks, achieving 59.7 PQ for multimodal panoptic and 78.2 mIoU for semantic segmentation, and also sets the new state of the art on DeLiVER. The source code is publicly available at: https://github.com/timbroed/CAFuser.
Index Terms:
Sensor fusion, semantic scene understanding, computer vision for transportation, deep learning for visual perception, multimodal semantic perception
@@ I Introduction
Extensive experiments show that our method, CAFuser (Condition-Aware Fuser), effectively learns to weigh different modalities in the fusion according to the present condition. CAFuser sets the new state of the art on DeLiVER [7] and ranks first on the public MUSES [3] benchmarks for both multimodal panoptic and semantic segmentation.
@@ III Method
We build on the recent OneFormer [17], using its head and segmentation framework as our starting point. In contrast to previous multimodal approaches, we introduce a single, shared backbone for all sensor modalities (see Fig. 2). By employing lightweight adapters for efficient feature transformation, this design significantly reduces model parameters while maintaining competitive performance. We initialize the backbone with ImageNet pre-training, which corresponds to the RGB camera modality. For pre-processing, we follow MUSES [3] and project each sensor’s data (e.g., lidar, radar) as a 3-channel image onto the RGB plane. Further, we normalize each modality over the entire dataset to ensure an input representation consistent with the RGB modality.
@@ III-A Multimodal Adapter
As the feature adapter, we employ a 2-layer MLP with a 4x reduction in hidden dimensions [5]. A learnable parameter $\alpha$ controls the weighting of adapted and original features. Each modality and each feature map from the Swin backbone’s 4-level feature pyramid is adapted using an individual adapter at each stage. For 4 modalities and 4 feature levels this results in 16 individual lightweight adapters. In Sec. IV-B we show that this setup allows us to reduce the parameters by 54% without any loss in performance.
@@ IV Experiments
We evaluate CAFuser on two multimodal driving datasets: MUSES [3] and DeLiVER [7]. MUSES contains 2500 real-world scenes with panoptic segmentation labels across eight conditions (day and night for rain, snow, fog, and clear weather). Each scene has RGB, lidar, radar, and event camera modalities, along with metadata describing the weather. We use the official SDK to project each modality onto the RGB plane. DeLiVER offers 7885 synthetic scenes for semantic segmentation in four adverse weather settings (cloudy, foggy, night, rainy), each with five corner-case artifacts (e.g. motion blur, lidar jitter). Each scene has RGB, projected lidar, depth, and event camera modalities. We generate a textual condition prompt similar to Sec. III-B with the template: “A synthetic [condition] driving scene with [case] artifacts.”.
CAFuser uses a Swin-T backbone [40] and follows the standard OneFormer training setup [17] with a batch size of 8. For MUSES, we train for 960 epochs with a 20% random drop of each modality [3] and for DeLiVER, we use 200k iterations, selecting the best checkpoint on the validation set.
@@ IV-A Comparison to The State of The Art
In Table I, we compare our model with existing panoptic segmentation methods, both using RGB-camera-only inputs and multimodal inputs. For RGB-camera-only panoptic segmentation, we compare to strong baselines such as Mask2Former [14], MaskDINO [16], and OneFormer [17]. Among these methods, OneFormer achieves the highest performance, highlighting the strength of this architecture. In a multimodal setting, MUSES [3] achieves the highest PQ with 53.9%. However, our method, which introduces CAF, achieves a new state-of-the-art result with a PQ score of 59.7%. The performance gain is especially notable in the night split with +7.6% margin over the second best method, compared to +1.9% in the day split. Since each time-of-day split includes all weather conditions (clear, fog, rain, and snow), these findings underscore the importance of adapting sensor weights when the RGB modality is less reliable.
As multimodal panoptic segmentation is still an emerging field, we further benchmark our model against the state of the art in multimodal semantic segmentation on both MUSES in Table II and DeLiVER in Table III. Given the OneFormer head simultaneously solves panoptic and semantic segmentation, we obtain one CAFuser model solving both tasks on MUSES. The results show that our model outperforms top methods such as CMNeXt [7] and GeminiFusion [8]. While these multimodal methods do improve over the RGB-camera-only methods, our CAFuser model significantly outperforms all multimodal and RGB-only methods and sets the new SOTA. This validates the strength of CAF, which allows optimal sensor integration under diverse environmental conditions and thus superior performance both in panoptic and semantic segmentation.
[…]
Method Clear Fog Rain Snow Day Night All
Mask2Former [14] 48.8 46.5 45.4 45.1 49.4 39.4 46.9
MaskDINO [16] 54.1 46.2 46.23 48.54 51.9 42.7 49.4
OneFormer [17] 58.3 53.7 53.4 53.8 57.6 47.8 55.2
HRFuser [6] 47.0 43.6 42.7 40.6 44.6 40.0 43.9
MUSES [3] 55.3 50.3 53.8 50.5 54.1 49.7 53.6
CAFuser-CAA 61.2 56.4 59.4 57.9 59.9 56.2 59.4
CAFuser-CA2 (Ours) 61.4 57.5 59.6 57.2 59.5 57.3 59.7
@@
TABLE III: Comparison of semantic segmentation methods on DeLiVER. C: RGB Camera, D: Depth, E: Events, L: Lidar
Method Modalities Backbone mIoU-val $\uparrow$ mIoU-test $\uparrow$
CMNeXt [7] CLDE MiT-B2 66.3 53.0
StitchFusion [28] CLDE MiT-B2 68.2 53.4
GeminiFusion [8] CLDE MiT-B2 66.9 54.5
CAFuser-CAA CLDE Swin-T 68.6 55.2
CAFuser-CA2 (Ours) CLDE Swin-T 67.8 55.6
@@ IV-B Ablation Studies
We perform all our ablations on MUSES due to its diverse real-world conditions and high-quality panoptic annotations.
Feature Adapter: We ablate the effect of our proposed feature adapter and CAF module in Tab. IV where we first create a strong baseline (row 2) by training OneFormer with the 4 parallel backbones proposed by MUSES. Using a shared backbone (row 3) significantly reduces the number of parameters (-55%), but also results in a significant performance drop (-0.9% PQ). Adding our proposed feature adapter (row 4) gains back all the lost performance (+0.9% PQ) while still having less than half of the original parameters. Adding our CAF module (row 5) gains another 0.4% in PQ, surpassing the performance of the larger, baseline model.
TABLE IV: Ablation study of our proposed modules on MUSES. Using a shared backbone reduces the parameters by 55% while the adapter and CAF module increase the PQ to a new SOTA.
@@
CAA Fusion: We conduct an ablation study to evaluate the impact of our CAA fusion method. As shown in Table V, we compare four fusion strategies: mean, random weights, learned weights, and CAA Fusion, which dynamically predicts modality weights using the CT. For the basic method that takes a simple mean over all modalities, we can see a large improvement of +1% PQ when adding the adapter (row 2) to the no-adapter baseline, highlighting the benefit of adapting features before fusion. Random fusion, where weights are randomly assigned to each modality, underperformed with a PQ of 58.5%. Learning static weights for each modality, without considering environmental conditions, reaches a similar performance to the mean fusion. Finally, our proposed CAA Fusion, which dynamically adjusts weights based on the CT, outperforms all other strategies with a PQ of 59.4%, demonstrating the effectiveness of adapting the fusion process to environmental conditions.
TABLE V: Ablation study using all four modalities on different weighted fusion strategies for our CAA module on MUSES.
@@
CA2 Fusion: In Tab. VI, we investigate the detailed design of our CA2 fusion and how to best utilize the CT for CAF. Appending the CT to the secondary modalities’ keys and values yields a small decrease in performance. Since the CT is derived from RGB features, mixing it with other modalities could introduce confusion to the fusion process.
In contrast, our proposed approach (row 3), where the CT is only appended to the RGB queries, achieves the highest PQ score of 59.7%. This design enables dynamic cross-attention tailored to environmental conditions, effectively guiding the fusion process and aligning with our intuition described in Sec. III-B and motivating our design choice.
TABLE VI: Ablation study on CA2 fusion design on MUSES. The Condition Token is appended to the RGB queries (Q) or the secondary modalities’ keys and values (K&V).
@@
Condition Loss: In Table VII, we assess the impact of our contrastive condition loss, by training an identical network, with and without applying the loss on the CT. As this loss is designed to ensure the CT effectively captures environmental conditions, we want to investigate if the model could learn this internally without additional supervision. Adding the condition loss improves the performance by +0.4% in PQ. This demonstrates that our contrastive loss is essential as an additional supervision to efficiently guide CAF.
TABLE VII: Ablation study on the guidance of the verbo-visual contrastive loss between $Q_{text}$ and the CT on MUSES.
@@
Figure 6: Qualitative panoptic segmentation results on MUSES with visualization of the four input modalities. Best viewed on a screen at full zoom.
Condition Encoding: In Table VIII, we ablate the use of different condition prompt encodings in the CT generation. Using a high-level condition description, such as simply classifying weather with a single token as e.g. clear or foggy, yields the same performance as without CAF (PQ 59.2%). However, our proposed detailed condition prompt construction, which captures nuanced aspects of the environment (e.g., distinguishing between snow on the ground versus snow in the air), results in a significant improvement of +0.5% in PQ, achieving a PQ of 59.7%. This suggests that detailed condition prompts provide more fine-grained information for the fusion process, enabling more effective sensor integration based on the actual environmental context.
Modalities: Table IX shows the performance of our model with different combinations of input modalities. Starting with RGB-only input (55.7% PQ), we observe consistent performance gains with each added modality: +3.0% PQ for lidar, +0.6% PQ for radar, and an additional +0.4% PQ for event camera, reaching 59.7% PQ. These results demonstrate that our approach effectively leverages all modalities, with each additional sensor contributing to performance. The flexibility of our design allows to generalize well across various input combinations.
TABLE VIII: Ablation study on the condition prompt encodings on MUSES. 1: simple high-level condition prompt (e.g. [“clear”]) 2: detailed condition prompts according to Sec. III-B.
@@
TABLE IX: Ablation for CAFuser with different input modalities on MUSES. Each added modality improves the PQ.
@@ V Conclusion
In this work, we introduced CAFuser, a novel condition-aware multimodal fusion framework for robust semantic perception in autonomous driving. By employing a shared backbone and modality-specific feature adapters, CAFuser efficiently aligns diverse sensor inputs into a common latent space while significantly reducing the model complexity. Our attention-based condition-aware fusion module dynamically adapts to environmental conditions guided by a Condition Token that is learned from the RGB input. This dynamic fusion enhances robustness and accuracy under challenging weather scenarios. We demonstrated that our method outperforms competing approaches both in panoptic and semantic segmentation, setting the new state of the art on MUSES and DeLiVER. Extensive ablation studies demonstrate the effectiveness of our proposed modules. These advancements make CAFuser a promising solution for enhancing semantic perception in autonomous driving and robotics, particularly under adverse environmental conditions.